HUBBLE SPACE TELESCOPE DAILY REPORT #2228 PERIOD COVERED: 0000Z (UTC) 10/13/98 - 0000Z (UTC) 10/14/98 Daily Status Report as of 287/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed NIC/1/2/3 7239 (Surface Compositions of Kuiper Belt Objects) The Near Infrared Camera (Cameras One, Two, and Three) was used to make observations of the Kuiper Belt object 1996TL66-2 in order to study its surface composition. The surfaces of Kuiper Belt Objects contain a record of formative and dynamical processes in the outer solar system. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 7618 (WF/PC-2 Cycle 7 Photometric Monitor UV/OPT STD) The WF/PC-2 was used to make observations of the standard star GRW+70D582 as part of a Cycle 7 photometric monitoring program. This ultraviolet spectrophotometric standard star was observed in a variety of filters and cameras to monitor the photometric stability and quantum efficiency of WF/PC-2 from the far-ultraviolet to near-infrared. The observations were completed as scheduled, and no problems were reported. 1.5 Completed Eight Sets of NIC/1/2/3 7963 (DARKs: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. As described in 2.1 and HSTAR 6713, the acquisition for the seventh iteration of this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as planned, and no further anomalies were noted. 1.6 Completed WF/PC-2 7626 (UV Throughput) The WF/PC-2 was used to make observations of PPL-15 as part of as ultraviolet throughput monitoring program. Observations were made of standard object GRW+70D582. The observations were completed as planned, and no problems were reported. 1.7 Completed STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.8 Completed Four Sets of NIC/2/3 8082 (NICMOS pure parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. As described in 2.1 and HSTAR 6713, the acquisition for the final two iterations of this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.9 Completed WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.10 Completed NIC/1/2/3 7876 (The Giant Branch Luminosity Function of M-31's Bulge) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the central bulge of the galaxy M-31 to study its giant branch luminosity function. A delineation of the differences in the luminous M giant population of the bulge of M-31 is of considerable interest for stellar synthesis work and for understanding the evolution of spiral bulges in general. The observations were completed as planned, and no problems were reported. 1.11 Completed FGS 7490 (Triton's Distorted Atmosphere) The Fine Guidance Sensors were used to make astrometric observations of Neptune's largest moon Triton in order to study the moon's distorted atmosphere. Observations were also made on several reference stars. The asymmetry of a stellar-occultation light curve for Triton observed in 1995 can be readily explained if Triton's middle atmosphere is distorted from spherical symmetry. The observations were executed as scheduled, and no anomalies were reported. 1.12 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.13 Completed NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star HD173740. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.15 Completed WF/PC-2 7427 (Saturn's Rings) The WF/PC-2 was used to make imaging observations of both the east and west apsides of Saturn's rings. Saturn's rings are a unique dynamical laboratory and continue to present challenging problems which require additional data to attack. The observations were completed as planned, and no problems were reported. 1.16 Completed NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around nearby, young main sequence stars. Because of the extreme youth of these objects, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed WF/PC-2/NIC/1/2/STIS/MA2 7561 (STIS GTO Observations) The WF/PC-2, Near Infrared Camera (Cameras One and Two) and the Space Telescope Imaging Spectrograph (MA2) were used to make observations of NGC-4194. The long slit capabilities of STIS will be used to investigate the stellar population and interstellar environment of the galaxy NGC-4194. Imaging of galaxies with WF/PC-2 has indicated the presence of knots of blue stars in many galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.18 Completed NIC/1/2/3 7847 (Near-IR Photometry of Candidate Companion to Proxima Centauri) The Near Infrared Camera (CameraS One, Two, and Three) was used to make observations of star Prox-Cen. This is to help study the star Proxima Centtauri which may have a low luminosity companion. The observations were completed as planned, and no anomalies were reported. 1.19 Completed STIS/MA1 7298 (A Snapshot Survey for UV bright Quasars at redshift below 3) The Space Telescope Imaging Spectrograph (MA1) was used to make ultraviolet snapshot observations of the bright quasar PKS0329-25 which has a redshift (Z) below 3.0. The goal is to find quasars bright enough in the ultraviolet for spectroscopic follow-up in the range of 900-1200 Angstroms. A wave length calibration image was also taken. The observation were completed as scheduled, and no anomalies were reported. 1.20 Completed STIS/CCD/MA1/MA2 7269 (The 1997/98 Eclipse of VV Cephei) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make ultraviolet spectrographic observations of the long-period eclipsing binary star HD208816 (a.k.a., VV-Cephei). This eclipse will provide an unprecedented opportunity to obtain spatially resolved information about the density, ionization and velocity structure in the outer atmosphere of the M supergiant primary. A wave length calibration observation was also taken. As described in 2.1 and HSTAR 6713, the acquisition for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as scheduled, and no further anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 14 Successful : 14 The acquisition at 286/190049Z defaulted to fine lock backup on FGS-3 when FGS-1 exceeded the scan step limit. HSTAR 6713 was written. The observations described in 1.5, 1.8, and 1.20 may have been affected. HSTAR 6714 described the acquisition at 287/030617Z which defaulted to fine lock backup on FGS-1 only when FGS-3 exceeded the scan step limit. Several proposals (WF/PC-2 7619 and NICMOS 7948) detailed in tomorrow's report may have been affected. Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 34 Successful : 34 2.3 Operations Notes: The two-week-long Hubble Deep Field - South campaign completed successfully on October 10. No significant problems were encountered, and data reduction is well underway. Several flanking field observations will take place as late as October 29. Raw and reduced data will be released on November 23. Further information may be obtained at: http://www.stsci.edu/ftp/science/hdf/hdfsouth/hdfs.html. The SSR EDAC error counter was reset four times per ROP SR-1. Six operational limits for the NICMOS instrument were modified per an operations request at 286/1833Z. As documented in HSTARs 6712 and 6715, NICMOS EDAC single bit errors occurred at 286/150155Z and at 287/070655Z. At 286/1846Z and at 287/0819Z the NICMOS internal log was dumped per an operations request. After the each occurrence, the ground system limit was modified, the final time to 25. Using ROP SR-3, the SSR science pointer was set twice for two 1 Meg realtime science data transmissions. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH